log base4(7x-8)=log base4(5x+3)
Answer by lwsshak3(6460)

log base4(7x-8)=log base4(5x+3)
log4(7x-8)-log4(5x+3)=0
place under single log
log[(7x-8)/(5x+3)]=0
convert to exponential form: base(4) raised to log of number(0)=number[(7x-8)/(5x+3)]
4^0=[(7x-8)/(5x+3)]=1
7x-8=5x+3
2x=11
x=5.5